OpenStudy (anonymous):

(D^4+6D^3+15D^2+20D+12)Y=0 higher order differential equation

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

you can try y=e^(rx) and solve the characteristic equation

OpenStudy (dan815):

Power series always works too

OpenStudy (anonymous):

i was trying to take common out like this 

OpenStudy (dan815):

uhhh

OpenStudy (dan815):

whats up with the 32

OpenStudy (dan815):

I think you should solve the characteristic equation

OpenStudy (dan815):

(D^4+6D^3+15D^2+20D+12)Y=0 Y=e^(rx) e^(rx)(r^4+6*r^3+15*r^2+20r+12)=0 (r^4+6*r^3+15*r^2+20r+12)=0 solve for r

zepdrix (zepdrix):

Is it a 32 or a 12? 0_o now I'm confused also

OpenStudy (dan815):

i think u added the 20 + 12 for some reason :P

OpenStudy (anonymous):

yep i m trying i assumed r=-2 but not getting the answer even with -1 so that means r = 0

OpenStudy (dan815):

r=0 wont work since we have a constant in the characteristic equation

OpenStudy (dan815):

(r^4+6*r^3+15*r^2+20r+12)=0 try -4 and -3

OpenStudy (dan815):

-6 too

OpenStudy (anonymous):

-4 IS WORKING

OpenStudy (dan815):

okay now use synthetic division and carry on

OpenStudy (anonymous):

but in my book the first power is -2 why ?

OpenStudy (dan815):

then -2 works too

OpenStudy (dan815):

yeah.. i dunno you need to try these numbers out properly and factor it

OpenStudy (anonymous):

wait let me try

OpenStudy (anonymous):

Now i have D^2 +2D+3

OpenStudy (anonymous):

@dan815

zepdrix (zepdrix):

So you had a repeated root at -2 I guess? :)

zepdrix (zepdrix):

\[\Large\rm D=\frac{-2\pm\sqrt{4-4(3)}}{2}\]I guess the last two roots are going to be complex, ya?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i got 4 roots -2,-2,-3,-1

OpenStudy (anonymous):

are they correct

zepdrix (zepdrix):

-3 and -1 would give you -4D for the middle term, not 2D. Those don't quite work out :(

zepdrix (zepdrix):

No more real roots after your -2 and -2. Have to use Quadratic Formula from here.

OpenStudy (anonymous):

i used factorization

zepdrix (zepdrix):

\[\Large\rm D=\frac{-2\pm\sqrt{-8}}{2}\]\[\Large\rm D=\frac{-2\pm2\sqrt{-2}}{2}\]\[\Large\rm D=-1\pm \sqrt2~i\]Ya these last two are complex roots :o Remember what your y solutions will look like when you get complex roots to your characteristic? :)

zepdrix (zepdrix):

..? :\

OpenStudy (anonymous):

oops sorry thanks for cleareing up the confusion

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

in scientific calculator do we have a direct way solving Quadratic equation in similar way

zepdrix (zepdrix):

Actually yes! If you graph it, y=x^4+6x^3+15x^2+20x+12 You'll get a parabola shape like this. https://www.desmos.com/calculator/9da3yxebbe When looking for roots, you're really looking for where the line touches the x-axis.